PHYSICS Q:

Without glasses a man can see objects clearly if they are no closer than 1.8 meters from his eyes. He walks into London Drugs and tries out some glasses there that have a power of +2.3 diopters. If he wears these glasses 1cm from his eyes, what is the shortest distance from his eyes at which he can see objects clearly?
Answer #1

Let us assume the shortest distance is d from the lens,

using lens formula,

1/-u + 1/v = 1/f = P

1/d + 1/-(1.80-0.01) = 2.3

    1/d - 1/1.79 = 2.3

d =0.35 m = 35 cm

shortest distance from eyes = 35 cm + 1 cm = 36 cm answer